What would be your specifications for the plane diffraction grating to view the mercury spectral line 577nm of the third order? Explain
1
Expert's answer
2020-06-08T10:09:48-0400

As per the question,

A large number of narrow rectangular slits having width in the order of the wave length of light and arranged side by side with equal opaque spaces is called a plane transmission diffraction grating. It is an extension of diffraction due to double slit.

wavelength of wave is given "(\lambda)=577nm"

m=3

We know that "d\sin \theta=m\lambda"

"\Rightarrow d\sin \theta=577\times 10^{-9}\times 3m"

"\Rightarrow d\sin\theta=1731\times 10^{-9}m"

"d\sin \theta=1.731\times 10^{-6}m"

"d\sin \theta=1.731\mu m"